OpenStudy (anonymous):

Anyone help me with this problem? Each hour, the hour hand of a clock moves This is 1/12 'th of the way around the circle. This is 1/12· 360° = 30°. Each minute, the hour hand moves This is 1/60'th of the distance it moves in one hour. This is 1/60· 30° = 0.5° Each minute, the minute hand moves 1/60'th of the way around the circle. This is 1/60· 360° = 6°. The angle is...?

OpenStudy (anonymous):

you didn't post a question yet

OpenStudy (anonymous):

the angle made by the difference between where the hand was and now is one unit of time later.

OpenStudy (anonymous):

@eseidl, I did. it's in the question

OpenStudy (anonymous):

How do I solve?

OpenStudy (anonymous):

Do I just multiply then add the numbers all up?

OpenStudy (anonymous):

What are you trying to solve for?

OpenStudy (anonymous):

Use the following information to find the angle between the hands of a clock at 5:15.

OpenStudy (anonymous):

so the degree =?

OpenStudy (anonymous):

Find the angle of the hour hand first: 5 is 5/12*360 around the clock, 5.15 is 1/4 *1/12 more around the clock. 5/12*360 = 150 1/4*1/12*360 = 7.5 So, the hour hand is 157.5 from 12, and the minute hand at 90 from 12. 157.5-90 = 67.5 (difference between the hour and minute hand) 157.5 (difference between hour and second (assuming the seconds are at 00)) 90 (difference between minute and second) Does that help?

OpenStudy (anonymous):

how do I solve?

OpenStudy (anonymous):

Um, so do I just add those all together?

OpenStudy (anonymous):

so is the answer 90 degrees?

OpenStudy (anonymous):

Your question: Use the following information to find the angle between the hands of a clock at 5:15. There are 3 different hands on the clock, could be seen as 3 points; A, B, C. You want to measure the degree at the origin, where they connect O. So there are 3 angles to find. AOB AOC BOC Where A is the Hour hand, B is the Minute hand and C is the Second hand and the time is 5.15.00: AOB = 157.5-90 = 67.5 AOC = 157.5 BOC = 90